From the probability mass function f(x) =2x+1/25



for x = 0,1,2,3,4, find the



cumulative distribution function of the random variable X. Using F(x), determine



the following probabilities.



(a) P(X ≤ 1)



(b)P(2 ≤ X < 4)



(c) P(X < 3)

1
Expert's answer
2022-03-08T04:32:03-0500

"F(x)=0" , for "x<0"

"F(x)={\frac {2*0+1} {25}}={\frac 1 {25}}" , for "0\u2264x<1"

"F(x)={\frac 1 {25}}+{\frac {2*1+1} {25}}={\frac 4 {25}}" , for "1\u2264x<2"

"F(x)={\frac 4 {25}}+{\frac {2*2+1} {25}}={\frac 9 {25}}" , for "2\u2264x<3"

"F(x)={\frac 9 {25}}+{\frac {2*3+1} {25}}={\frac {16} {25}}" , for "3\u2264x<4"

"F(x)={\frac {16} {25}}+{\frac {2*4+1} {25}}={\frac {25} {25}}=1" , for "x\u22654"

a) "P(X≤1)=F(1)={\frac 4 {25}}"

b) "P(2≤X<4)=F(3)-F(1)={\frac {16} {25}} - {\frac 4 {25}}={\frac {12} {25}}"

c) "P(X<3)=F(2)={\frac 9 {25}}"